OpenStudy (driftracer305):

the reflecting dish of a parabolic microphone has a cross-section in the shape of a parabola. the microphone itself is placed on the focus of the parabola. if the parabola is 24 inches wide and 9 inches deep, how far from the vertex should the microphone be placed? 6 in , 8 in, 16 in, , 4 in

OpenStudy (driftracer305):

@ganeshie8

OpenStudy (driftracer305):

@phi @campbell_st

OpenStudy (mackenzie2013):

4 inches.

OpenStudy (mackenzie2013):

Place the vertex at the origin. Then the parabola has equation 4py = x² 4p(9) = 144 p = 4 in

OpenStudy (mackenzie2013):

I have no comment..

OpenStudy (mackenzie2013):

My lips are sealed. Looks like you're the one who googled it :P

OpenStudy (driftracer305):

lol........... Ms. Smarty @Mackenzie2013 .......... good one..!! haha

OpenStudy (mackenzie2013):

Thank ya! haha

OpenStudy (driftracer305):

uhh @campbell_st ....... pls help?!??

OpenStudy (campbell_st):

the solution you have been given is correct... here is a diagram showing why...  so the general form I use is \[x^2 = 4ay\] where a is the focal length... or distance from the vertex to the focus... so substituting x = 12 and y = 9 it becomes \[144 = 4 \times a \times 9\] solving for a will give the focal length and subsequently how far above the vertex the focus is. hope it makes sense.